OpenStudy (anonymous):

a married couple, purchased their home in December 1989 for $89,000. The house was appraised in January 2008 for $325,000. What was the annual rate of appreciation on the home during this period?

OpenStudy (amistre64):

compounded yearly eh...

OpenStudy (amistre64):

325 = 89(1+r)^t right?

OpenStudy (amistre64):

3.652 = (1+r)^18 if im doing it right :)

OpenStudy (anonymous):

ok how do you get the r in the bracket...this is where I am confused

OpenStudy (amistre64):

the r = rate of appreciation correct? if we take the initial value, and apply the "interest" to it we get: 89 + 89(rate); which factors to: 89(1+ rate) the rate is compunded every year for a total of 18 years. 89(1+rate)^18 = 325

OpenStudy (anonymous):

Is the t = 19?

OpenStudy (anonymous):

yes the t=19

OpenStudy (amistre64):

log(3.652) = log[(1+r)^18] log(3.652) = 18 log(1+r) 3.652^(1/18) = 1+r 3.652^(1/18) -1 = r

OpenStudy (amistre64):

dec 89 is basically jan 90

OpenStudy (amistre64):

8+10 = 2008

OpenStudy (amistre64):

I get about 7.5%

OpenStudy (anonymous):

ok... well the answers that I have to choose from are: 6.69% 6.26% 6.87% 7.05%

OpenStudy (amistre64):

im close then ;)

OpenStudy (amistre64):

work it backwards to see what fits, thats what I do with multiple choices

OpenStudy (anonymous):

If you use 19 then you get 7.05%

OpenStudy (anonymous):

lol...yes...but I am trying to find out how to come up with that on a BAII Plus calculator. And log is not on this one

OpenStudy (amistre64):

89(1.0705)^18 use the calculator on the computer.... programs, accessories, calculator usually

OpenStudy (anonymous):

ok....we are only allowed to use our calculators in this exam, so I have to figure out how to do it on this one. Stupid Business Math! LOL

OpenStudy (amistre64):

324.7407 if you use ^19 ;)

OpenStudy (amistre64):

without a log button...you got exponent button?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

can you key in a random exponent? like 1/19?

OpenStudy (amistre64):

[325/89]^(1/19) -1

OpenStudy (amistre64):

0.070544965363919914409636326010341

OpenStudy (anonymous):

I use the button on the calculator that has a radical sign with the x above the radical sign and a y under the radical sign Type in 3.652 then that sign and then 19 =

OpenStudy (amistre64):

1/19 not 19 :) way different value for that lol

OpenStudy (amistre64):

ohh....radical sign, gotcha

OpenStudy (anonymous):

wow...mind spinning....lol

OpenStudy (amistre64):

the answers 7.05 ;) dont lose sight of that one lol

OpenStudy (anonymous):

yes....thanks for coming up with the appropriate answer...but I am trying to find out how to punch in this formula. I am a math idiot.... wish I wasn't

OpenStudy (anonymous):

OMG......figured it out....yipee. Sooooo love you guys!!!

OpenStudy (amistre64):

take a pic of the calculator and post it :)

OpenStudy (amistre64):

or not lol :)

OpenStudy (amistre64):

yay!!! youre welcome :)

OpenStudy (anonymous):

ok....will do

OpenStudy (anonymous):

wish I knew how to use this dang thing....or that I knew how to do math

OpenStudy (amistre64):

your math might be fine, that calculator looks mean tho lol

OpenStudy (anonymous):

it is....we are all used to a regular scientific one. So we don't know how to use all the buttons that would probably make it much simpiler

OpenStudy (amistre64):

heres my other one ;)

OpenStudy (amistre64):

it didnt post..... how odd

OpenStudy (amistre64):

try it again

OpenStudy (anonymous):

NOW THAT'S MY KIND OF CALCULATOR! Bahahahaha

OpenStudy (amistre64):

:) LOL

OpenStudy (anonymous):

well you ARE now my Superhero!! Thanks so much. And I will definitely be back, I assure you! LOL

OpenStudy (amistre64):

have a great time ;)